OpenStudy (anonymous):

How many solutions exist for the system, x + y = 1 and y = -x + 1? A. no solution B. one solution C. infinitely many solutions I think its C.... am I right?

OpenStudy (anonymous):

thnk u for ur answer Salv.

OpenStudy (anonymous):

it is C

OpenStudy (anonymous):

mmm i think it is c

OpenStudy (anonymous):

sorry, don't know what happened to my post

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

it's C: if you tried substituting either x=-y+1 or y=-x+1 into x+y=1 you would find that the variable will cancel leaving you with a consistent equation 1=1. The fact that the variable cancels means that it may be anything. Since this is true for both x and y (which ever substitution you choose to make) then there are infinitely many solutions.